OpenStudy (anonymous):

how to find the derivative of e^xy = 2x+y

OpenStudy (anonymous):

implicit differentiation for this one use the chain rule and product rule

OpenStudy (anonymous):

\[e^{xy}\times \left(y+xy'\right)=2+y'\] solve for \(y'\)

OpenStudy (anonymous):

could u explain how to got y +xy'

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are taking the derivative of \(e^{xy}\) with respect to \(x\) the derivative of \(e^{whatever}\) is \(e^{whatever}\times (whatever)'\)so your real job is to take the derivative of \(xy\) with respect to \(x\) for this you need the product rule

OpenStudy (anonymous):

we are thinking of \(y\) as a function of \(x\) so it is like taking the derivative of \[xf(x)\] by the product rule you get \[f(x)+xf'(x)\] but we are writing \(y\) instead of \(f(x)\) and \(y'\) instead of \(f'(x)\) write instead \[y+xy'\]

OpenStudy (anonymous):

ok. Would you mind helping with another problem ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i'm trying to find the derivative of e^x/ 1-e^2x , i did it, but I get a different answer than the book

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=e^x, f'(x)=e^x, g(x)=1-e^{2x}, g'(x)=-2e^{2x}\]

OpenStudy (anonymous):

before any simplification you get \[\frac{(1-e^{2x})e^x-e^x\times (-2e^{2x})}{(1-e^{2x})^2}\]

OpenStudy (anonymous):

can simplify this a bit for sure

OpenStudy (anonymous):

yes that is what I have. Can you simplify it, bc when I do it i get a wrong answer

OpenStudy (anonymous):

\[\frac{(1-e^{2x})e^x-e^x\times (-2e^{2x})}{(1-e^{2x})^2}\] \[\frac{(1-e^{2x})e^x+e^x\times (2e^{2x})}{(1-e^{2x})^2}\]

OpenStudy (anonymous):

\[\frac{e^x(1-e^{2x}+2e^{2x})}{(1-e^{2x})^2}\]

OpenStudy (anonymous):

wait how did you get -2e^2x before simplifying

OpenStudy (anonymous):

\[\frac{e^x(1+e^{2x})}{(1-e^{2x})^2}\]

OpenStudy (anonymous):

wait, I know

OpenStudy (anonymous):

I just missed a negative when I was taking the derivative of the bottom

OpenStudy (anonymous):

ahh i see

OpenStudy (anonymous):

Sorry to ask, but could you help me with another question